# Average Cost Optimality of Partially Observed MDPs: Contraction of Non-Linear Filters and Existence of Optimal Solutions

### 21 Dec 2023 v1

@@ 3. Concluding Remarks
The average cost optimality is a challenging problem, especially under the average cost criterion. In this paper, we presented explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. A comparison with the related literature and several examples are presented. Notably, our paper appears to be the first one where a contraction result is presented and utilized for the optimality analysis. Further research will involve providing implications of our contraction based analysis on approximations and learning theoretic results.
4. Appendix
4.1. Proof sketch of Theorem 2.3
@@ 4. Appendix
(14) $=\sup_{f\in\operatorname{Lip}(\mathcal{Z}),\;\left\lVert f\right\rVert_{\infty}\leq D/2}\bigg{|}\int_{\mathbb{Y}}f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)P\left(dy_{1}\mid z_{0}^{\prime},u\right)$
@@
As a result, we get $\tilde{g}/\alpha\in Lip(\mathbb{X})$.
@@
where $w_{y_{1}}=\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)-z_{1}\left(z_{0},u,y_{1}\right)\right)$ which is a signed measure on $\mathbb{X}$. By the Measurable Selection Theorem choose measurable $g_{y}\in\arg\sup_{g\in\operatorname{Lip}(\mathbb{X})}\left(\int_{\mathbb{X}}g(x)w_{y}(dx)\right).$
After that we can continue with the equation (4.1),
$\int_{\mathbb{Y}}\sup_{g\in\operatorname{Lip}(\mathbb{X})}\left(\int_{\mathbb{X}}g(x_{1})w_{y_{1}}(dx_{1})\right)P\left(dy_{1}\mid z_{0},u\right)$
@@
If we take the supremum of the equation over all $f\in Lip(\mathcal{Z})$, then by using the inequalities (17),(22), and (4.1), we can write:

### 13 Mar 2024 v2

@@ 4. Concluding Remarks
The average cost optimality is a challenging problem, especially under the average cost criterion. In this paper, we presented explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. A comparison with the related literature and several examples are presented. Notably, our paper appears to be the first one where a contraction result is presented and utilized for the optimality analysis. We finally presented several implications of our analysis and existence result on approximations, learning, and robustness.
Appendix A Proof of Theorem 2.2
@@
(23) $=\sup_{f\in\operatorname{Lip}(\mathcal{Z},1),\;\left\lVert f\right\rVert_{\infty}\leq D/2}\bigg{|}\int_{\mathbb{Y}}f\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)\right)P\left(dy_{1}\mid z_{0}^{\prime},u\right)$
@@
As a result, we get $\tilde{g}/\alpha\in Lip(\mathbb{X},1)$.
@@
where $w_{y_{1}}=\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)-z_{1}\left(z_{0},u,y_{1}\right)\right)$ which is a signed measure on $\mathbb{X}$. By the Measurable Selection Theorem choose measurable
$g_{y}\in\arg\sup_{g\in\operatorname{Lip}(\mathbb{X},1)}\left(\int_{\mathbb{X}}g(x)w_{y}(dx)\right).$
After that we can continue with the equation (A),
$\int_{\mathbb{Y}}\sup_{g\in\operatorname{Lip}(\mathbb{X},1)}\left(\int_{\mathbb{X}}g(x_{1})w_{y_{1}}(dx_{1})\right)P\left(dy_{1}\mid z_{0},u\right)$
@@
If we take the supremum of the equation over all $f\in Lip(\mathcal{Z})$, then by using the inequalities (26),(31), and (A), we can write: